Calculate the present value of the after tax net returns to land in the 7th year if thereal pre-tax net returns to land today are $250, real net returns to land are assumedto increase by 5% each year, inflation is 4%, the marginal tax rate is 20%, and thepretax risk adjusted discount rate is 10%.

a.

270.10

b.

216.08

c.

105.27

d.

163.84
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

Fn = Fo (1+g) n = 250(1.05)7 = 351.775

Nominal net returns = 351.775 (1.04)7 = 462.912

After tax return = 462.912 (1 - 0.2) = 370.33

After-tax, risk adjusted discount rate = .1(1-.2) = 0.08 ot 8%

PV(after-tax net return in 7th year = 370.33 (1+.08)-7 = 216.08